# MVGamba: Unify 3D Content Generation as State Space Sequence Modeling

### v2

@@ Abstract
Recent 3D large reconstruction models (LRMs) can generate high-quality 3D content in sub-seconds by integrating multi-view diffusion models with scalable multi-view reconstructors. Current works further leverage 3D Gaussian Splatting as 3D representation for improved visual quality and rendering efficiency. However, we observe that existing Gaussian reconstruction models often suffer from multi-view inconsistency and blurred textures. We attribute this to the compromise of multi-view information propagation in favor of adopting powerful yet computationally intensive architectures (e.g., Transformers). To address this issue, we introduce MVGamba, a general and lightweight Gaussian reconstruction model featuring a multi-view Gaussian reconstructor based on the RNN-like State Space Model (SSM). Our Gaussian reconstructor propagates causal context containing multi-view information for cross-view self-refinement while generating a long sequence of Gaussians for fine-detail modeling with linear complexity. With off-the-shelf multi-view diffusion models integrated, MVGamba unifies 3D generation tasks from a single image, sparse images, or text prompts. Extensive experiments demonstrate that MVGamba outperforms state-of-the-art baselines in all 3D content generation scenarios with approximately only 0.1×0.1$\times$0.1 × of the model size.
@@ 1 Introduction
More concretely, we first patchify the multi-view images into $N$ tokens and rearrange them according to a cross-scan order [27], resulting in $4\times N$ image tokens for selective scanning. These tokens are then processed through a series of Mamba blocks for state space sequence modeling. Subsequently, we feed the output Gaussian sequence into a lightweight Multi-Layer Perceptron (MLP) for channel-wise knowledge selection, followed by a set of linear decoders to obtain the Gaussian parameters representing high-quality 3D content (Sec. 3.2). Compared to previous LRMs [28, 11, 29], our MVGamba features many computationally efficient components: a single-layer 2D convolution image tokenizer replaces the pre-trained DINO [30] transformer encoder, a lightweight MLP combined with linear decoders replaces the deep MLP decoder, and most importantly, linear complexity Mamba blocks replace quadratic complexity Transformer blocks (Figure 2(b)). Together, these designs ensure efficient training and inference while achieving higher generation quality (Sec. 4). Moreover, to directly convert the generated Gaussians into smooth textured polygonal meshes, we alternatively incorporate a 3DGS variant — 2DGS [31] — for accurate geometric modeling and mesh extraction.
Figure 2: (a) Previous Gaussian reconstruction models sacrifice the integrity of multi-view information for computationally intensive architectures, resulting in multi-view inconsistency and blurred textures. (b) The computational complexity comparison: Transformer blocks have quadratic complexity $O(N^{2})$ [32], while Mamba blocks achieve linear complexity $O(N)$ [25].
@@ 3.2 SSM-based Gaussian reconstructor
In this way, we convert the rotation prediction into a $32$-class classification task in a fully differentiable way, which allows for direct selection via the $\operatorname{argmax}$ operation during inference. We refer to Appendix C for more detailed explanations. These decoded Gaussians are finally passed into the differentiable rasterization pipeline [21] for image-level supervision.
@@ 4 Experiment
Figure 4: Qualitative comparison in image-to-3D and text-to-3D generation. Please refer to Appendix B for more generation results.
@@ 4.1 Experimental Settings
Implementation details. MVGamba is trained on 32 NVIDIA A100 (80G) with batch size 512 for about 2 days. We adopt gradient checkpointing and mixed-precision training with BF16 data type to ensure efficient training and inference. We use the AdamW optimizer with learning rate $1\times 10^{-3}$ and weight decay $0.05$, following a linear learning rate warm-up for $15$ epochs with cosine decay to $1\times 10^{-5}$. The output Gaussians are rendered at $512\times 512$ resolution for mean square error loss and resized to $256\times 256$ for LPIPS loss for memory efficiency. The trade-off coefficients that balancing each loss were set as $\lambda_{\text{mask}}=1.0$, $\lambda_{\text{LPIPS}}=0.6$ and $\lambda_{\text{reg}}=0.001$. We also follow the common practice [19] to clip the gradient with a maximum norm of $1.0$. The detail of MVGamba model configuration is included in Appendix C.
@@ 4.2 Comparison against Baselines
Text-to-3D generation. Similar to the single image-to-3D task, we compare MVGamba with several optimization-based methods and feed-forward models LGM, TripoSR, and Triplane-Gaussian, using various prompts. For methods that only accept a single image as input, such as TripoSR and Triplane-Gaussian, we use DALL·E 3 [77] to transform the given text into a single image. Note that for more fair comparison, MVGamba and LGM take the same generated multi-view inputs. Figure 4 shows that MVGamba excels at generating plausible geometry and highly realistic texture. Due to its effective ray embedding and multi-view diffusion models, MVGamba is also free from the multi-face problems that frequently occur in optimization-based methods. We further randomly selected 50 prompts from the Dreamfusion [1] gallery and used CLIP Precision and CLIP scores [78, 79] to measure appearance quality and alignment with the given prompt. Our high generation quality is also vividly reflected in Table 1, where MVGamba consistently ranks the highest among all baselines.
Sparse-view reconstruction. Given the same sparse-view inputs, we compare MVGamba with SparseNeuS [80] (trained in One-2-3-45), SparseGS [81] and LGM [19] that are capable of generating 3D Gaussians. We visualize the reconstruction results of MVGamba in Figure 5 and compare against baselines on the GSO [76] dataset with 32 randomly selected views for each object. As shown in Figure 2, MVGamba is able to faithfully reproduce high-frequency details with accurate geometric modeling. Due to the limited space, the quantitative evaluation is presented in Appendix B.
@@ 6 Conclusion
The Appendix is organized as follows:
  • Appendix A: elaborates more details about our MVGamba pipeline. Specifically, we detailed the implementation of inference pipeline and mesh extraction.
  • Appendix B: showcases more qualitative and quantitative experiment results.
  • Appendix C: provides further details on model design and configuration.
  • Appendix D: discusses the limitations and potential future progress of MVGamba.
@@ Appendix A Implementation Details
Appendix B More Results
@@
We include more qualitative experiment on image-to-3D generation with rendering results in Figure 9 for novel view synthesis and normal map generation in Figure 10.
@@
We leave the qualitative comparison on sparse-view reconstruction task here. We compare the well-adpoted PSNR, SSIM, LPIPS for novel view synthesis and Chamfer Distance (CD) and Volume IoU (VIoU) for geometric evaluation. As shown in Table 3, MVGamba outperforms all baselines across all metrics, even though SparseNeuS require 4 times more input views, while maintaining fast inference and rendering speed.
@@
Figure 10: RGB images and normal maps generated by MVGamba-2DGS.
Appendix C Model Details
@@
Appendix D Limitations
Honestly, though MVGamba achieves promising results, it still has several limitations. (1) As we model the Gaussian sequence causally, MVGamba can sometimes fail if the depth of the front view is not estimated correctly (Figure 8 top). Fortunately, we empirically find that this limitation can be mitigated by manually changing the input order. For example, using a side-view as the first input allows our model to generate satisfactory 3D content, as the side-view contains sufficient depth information (Figure 8 bottom). In the future, we may explore automatic ways to optimize the input order to enhance robustness. (2) The generation quality of MVGamba is highly dependent on the four input views provided by off-the-shelf multi-view diffusion models [14, 13]. However, current multi-view diffusion models are far from perfect, known to exhibit 3D inconsistencies [19, 9, 84], and limited to a resolution of 256 $\times$ 256. We expect that our model’s performance can be seamlessly boosted with the advancements in multi-view diffusion models in future work.

### v3

@@ Abstract
Recent 3D large reconstruction models (LRMs) can generate high-quality 3D content in sub-seconds by integrating multi-view diffusion models with scalable multi-view reconstructors. Current works further leverage 3D Gaussian Splatting as 3D representation for improved visual quality and rendering efficiency. However, we observe that existing Gaussian reconstruction models often suffer from multi-view inconsistency and blurred textures. We attribute this to the compromise of multi-view information propagation in favor of adopting powerful yet computationally intensive architectures (e.g., Transformers). To address this issue, we introduce MVGamba, a general and lightweight Gaussian reconstruction model featuring a multi-view Gaussian reconstructor based on the RNN-like State Space Model (SSM). Our Gaussian reconstructor propagates causal context containing multi-view information for cross-view self-refinement while generating a long sequence of Gaussians for fine-detail modeling with linear complexity. With off-the-shelf multi-view diffusion models integrated, MVGamba unifies 3D generation tasks from a single image, sparse images, or text prompts. Extensive experiments demonstrate that MVGamba outperforms state-of-the-art baselines in all 3D content generation scenarios with approximately only 0.1×0.1$\times$0.1 × of the model size. The codes are available at https://github.com/SkyworkAI/MVGamba.
@@ 1 Introduction
More concretely, we first patchify the multi-view images into $N$ tokens and rearrange them according to a cross-scan order [27, 28], resulting in $4\times N$ image tokens for selective scanning. These tokens are then processed through a series of Mamba blocks for state space sequence modeling. Subsequently, we feed the output Gaussian sequence into a lightweight Multi-Layer Perceptron (MLP) for channel-wise knowledge selection, followed by a set of linear decoders to obtain the Gaussian parameters representing high-quality 3D content (Sec. 3.2). Compared to previous LRMs [29, 11, 30], our MVGamba features many computationally efficient components: a single-layer 2D convolution image tokenizer replaces the pre-trained DINO [31] transformer encoder, a lightweight MLP combined with linear decoders replaces the deep MLP decoder, and most importantly, linear complexity Mamba blocks replace quadratic complexity Transformer blocks (Figure 2(b)). Together, these designs ensure efficient training and inference while achieving higher generation quality (Sec. 4). Moreover, to directly convert the generated Gaussians into smooth textured polygonal meshes, we alternatively incorporate a 3DGS variant — 2DGS [32] — for accurate geometric modeling and mesh extraction.
Figure 2: (a) Previous Gaussian reconstruction models sacrifice the integrity of multi-view information for computationally intensive architectures, resulting in multi-view inconsistency and blurred textures. (b) Comparison of FLOPs between self-attention in Transformers and SSM in Mamba. Detailed FLOPs data are provided in Table 3.
@@ 3.2 SSM-based Gaussian reconstructor
In this way, we convert the rotation prediction into a $32$-class classification task in a fully differentiable way, which allows for direct selection via the $\operatorname{argmax}$ operation during inference. We refer to Appendix D for more detailed explanations. These decoded Gaussians are finally passed into the differentiable rasterization pipeline [21] for image-level supervision.
@@ 4 Experiment
Figure 4: Qualitative comparison in image-to-3D and text-to-3D generation. Please refer to Appendix C for more generation results.
@@ 4.1 Experimental Settings
Implementation details. MVGamba is trained on 32 NVIDIA A100 (80G) with batch size 512 for about 2 days. We adopt gradient checkpointing and mixed-precision training with BF16 data type to ensure efficient training and inference. We use the AdamW optimizer with learning rate $1\times 10^{-3}$ and weight decay $0.05$, following a linear learning rate warm-up for $15$ epochs with cosine decay to $1\times 10^{-5}$. The output Gaussians are rendered at $512\times 512$ resolution for mean square error loss and resized to $256\times 256$ for LPIPS loss for memory efficiency. The trade-off coefficients that balancing each loss were set as $\lambda_{\text{mask}}=1.0$, $\lambda_{\text{LPIPS}}=0.6$ and $\lambda_{\text{reg}}=0.001$. We also follow the common practice [19] to clip the gradient with a maximum norm of $1.0$. The detail of MVGamba model configuration is included in Appendix D.
@@ 4.2 Comparison against Baselines
Text-to-3D generation. Similar to the single image-to-3D task, we compare MVGamba with several optimization-based methods and feed-forward models LGM, TripoSR, and Triplane-Gaussian, using various prompts. For methods that only accept a single image as input, such as TripoSR and Triplane-Gaussian, we use DALL·E 3 [77] to transform the given text into a single image. Note that for more fair comparison, MVGamba and LGM take the same generated multi-view inputs. Figure 4 shows that MVGamba excels at generating plausible geometry and highly realistic texture. Due to its effective ray embedding and multi-view diffusion models, MVGamba is also free from the multi-face problems that frequently occur in optimization-based methods. We further randomly selected 50 prompts from the Dreamfusion [1] gallery and used CLIP Precision and CLIP scores [78, 79, 80] to measure appearance quality and alignment with the given prompt. Our high generation quality is also vividly reflected in Table 1, where MVGamba consistently ranks the highest among all baselines.
Sparse-view reconstruction. Given the same sparse-view inputs, we compare MVGamba with SparseNeuS [81] (trained in One-2-3-45), SparseGS [82] and LGM [19] that are capable of generating 3D Gaussians. We visualize the reconstruction results of MVGamba in Figure 5 and compare against baselines on the GSO [76] dataset with 32 randomly selected views for each object. As shown in Figure 2 and Figure 6, MVGamba is able to faithfully reproduce high-frequency details with accurate geometric modeling. Due to the limited space, the quantitative evaluation is presented in Appendix C.
Figure 6: RGB images and normal maps rendered by MVGamba-2DGS.
@@ 5 Ablation and Discussion
Q3: What is the limitation of MVGamba?
A3: Honestly, though MVGamba achieves promising results, it still has several limitations. (1) As we model the Gaussian sequence causally, MVGamba can sometimes fail if the depth of the front view is not estimated correctly (Figure 8 top). Fortunately, we empirically find that this limitation can be mitigated by manually changing the input order. For example, using a side-view as the first input allows our model to generate satisfactory 3D content, as the side-view contains sufficient depth information (Figure 8 bottom). In the future, we may explore automatic ways to optimize the input order to enhance robustness. (2) The generation quality of MVGamba is highly dependent on the four input views provided by off-the-shelf multi-view diffusion models [14, 13]. However, current multi-view diffusion models are far from perfect, known to exhibit 3D inconsistencies [19, 9, 83], and limited to a resolution of 256 $\times$ 256. We expect that our model’s performance can be seamlessly boosted with the advancements in multi-view diffusion models in future work.
@@ 6 Conclusion
7 Acknowledgments
This project is supported by Kunlun 2050 Research, Skywork AI and Agency for Science, Technology AND Research, and by the National Research Foundation, Singapore, under its Medium Sized Center for Advanced Robotics Technology Innovation. Pan Zhou was supported by the Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grants (project ID: 23-SIS-SMU-028 and 23-SIS-SMU-070).
@@ Appendix A Implementation Details
Appendix B More Discussion
B.1 Comparison with concurrent works.
We illustrate the two current sequence modeling paradigms for Gaussian LRMs: Compressed vs. Direct Sequence Modeling (ours) in Figure 9. Below, we elaborate on these two paradigms to highlight the unique approach of our proposed paradigm compared to recent transformer-based approaches, e.g., GRM, GS-LRM.
  • (a) Compressed Sequence Modeling: This approach tokenizes the input into a compact representation, processes it through a series of transformer blocks, and then upsamples to produce the 3DGS parameters. This paradigm is represented by the concurrent GS-LRM and GRM.
  • (b) Direct Sequence Modeling (ours): This approach tokenizes and expands the input into a sufficiently long sequence of tokens through cross-scan operations, which are then directly processed by a series of mamba blocks to generate the 3DGS parameters.
Our proposed paradigm (b) offers several significant benefits:
  • Efficient long sequence modeling: It accommodates larger spatial dimensions to capture and preserve fine-grained details while mitigating information loss typically caused by upsamplers, such as the zero-padding in de-convolution layers. This benefits accurate geometry and texture reconstruction. As demonstrated in Table 3, Mamba’s linear computational complexity allows for a favorable balance between computational cost and long-sequence modeling capacity, enabling efficient processing of high-resolution inputs without prohibitive memory requirements.
  • Easy of optimization: It directly models the 3DGS sequence without any upsampler, hence establishing a more straightforward relationship between 3DGS parameters (e.g., position, color) and the modeled sequence. This eases the non-convex optimization in inverse graphics scenario, as discussed in PixelSplat, and helps learn accurate geometry and textures.
  • Cross-view self-refinement: It efficiently incorporates multi-view information causally from the initial condition onward, thereby enabling the refinement of inconsistent parts based on earlier views and generated tokens.
Figure 9: The illustration of two type of Gaussian LRMs. (a) Compressed sequence modeling (b) Direct long-sequence modeling (ours).
Table 3: Theoretically calculated FLOPs comparison between self-attention in Transformer and SSM in Mamba. Here, dimension $D=512$; GFLOPS of both modules are calculated according to Equations (5) and (6) in Vision Mamba [26].
Model / Length 1024 2048 4096 8192 16384 32768
Self-attention / GFLOPs 2.15 6.44 21.47 77.31 292.06 1133.87
SSM / GFLOPs 0.07 0.13 0.27 0.54 1.07 2.15
B.2 The function of pixel-aligned Gaussian
We conducted an additional experiment using a non-pixel-aligned transformer model for 3DGS prediction. As discussed in Appendix B.1, transformer-based methods typically process a compressed sequence considering the computational budget. We then followed OpenLRM to model 4096 tokens, and adopt a de-convolution layer to up-sample the 3DGS tokens from 4096 to 16384, similar to TGS for intricate 3D modeling. However, as shown in Figure 11, the non-pixel-aligned transformer yielded inferior performance, with broken geometry and extremely blurred textures after $300$ epochs of training. We attribute this to the inherent non-convex optimization challenge in the inverse graphics scenario, as mentioned in pixelSplat [85], which may be further amplified by the upsampler. This preliminary result may also explain why pixel-aligned approaches are proposed and are adopted by recent transformer-based approaches with upsamplers. On the other hand, our proposed direct sequence modeling paradigm allows for more fine-grained detail modeling and exhibits a cross-view self-refinement ability. Moreover, our direct sequence modeling paradigm can also ease optimization, which provides a new feasible way for training Gaussian LRMs.
Figure 11: The generation results of non-pixel-aligned transformer-based Gaussian LRM, resulting in extremely broken geometry and blurred texture.
Appendix C More Results
@@
We include more qualitative experiment on image-to-3D generation with rendering results in Figure 10 for novel view synthesis and in Figure 12 for normal map generation. We also include a qualitative comparison using samples from the GRM paper in Figure 13. It indicates that MVGamba captures better geometries and details with significantly smaller model size. Note that we could not directly compare with GS-LRM as it only presents a single-view results in Figure 7 of their paper.
Figure 12: RGB images and normal maps rendered by MVGamba-2DGS.
Figure 13: More comparisons with concurrent state-of-the-art large Gaussian models. Note that the generation results for GRM were extracted from its paper and official project page. A direct comparison with GS-LRM is currently unfeasible, as it is not open-source, and only a single view is presented in Figure 7 of the GS-LRM paper.
@@
We leave the qualitative comparison on sparse-view reconstruction task here. We compare the well-adpoted PSNR, SSIM, LPIPS for novel view synthesis [86] and Chamfer Distance (CD) and Volume IoU (VIoU) for geometric evaluation. As shown in Table 4, MVGamba outperforms all baselines across all metrics, even though SparseNeuS require 4 times more input views, while maintaining fast inference and rendering speed.
@@
Appendix D Model Details
@@
Appendix E Licenses
Datasets:
  • Objaverse [7]: ODC-By v1.0 license
Pre-trained models:
  • ImageDream [13]: Apache-2.0 license
  • MVDream [14]: MIT License
  • SAM [84]: Apache-2.0 license